OpenStudy (anonymous):

which sentence is the most accurate definition of an identity? A) an identity is an equation that is true for all real numbers B) an identity is an equation that is true for all real values in its domain C) an identity in an equation that is true for all real values in its range D) an identity is an equation that can be verified by substituting input values.

OpenStudy (anonymous):

i know A isn't right because it's not always true

OpenStudy (anonymous):

Good, then it's c.

OpenStudy (anonymous):

it's not c

OpenStudy (anonymous):

Is it d? Amazing.

OpenStudy (anonymous):

How about doing the work yourself.

OpenStudy (anonymous):

that was a little mean.

OpenStudy (anonymous):

i would go with B since it has to be true, but would only make sense for numbers in the domain

OpenStudy (anonymous):

like for example \[\frac{1}{x}=\frac{2}{2x}\] is an identity ( a rather simple one) but is not true if \(x=0\)

OpenStudy (anonymous):

In mathematics an identity is an equality relation A = B, such that A and B contain some variables and A and B produce the same value as each other regardless of what values

OpenStudy (anonymous):

i think it is D

OpenStudy (anonymous):

i thought D was true as well because you would already have the values needed to check

OpenStudy (anonymous):

another county heard from you do not verify an identity by checking with values you do it mostly by algebra

OpenStudy (anonymous):

\[2(x+3)=2x+6\] is an identity via the distributive property, not by checking with numbers

OpenStudy (anonymous):

Boo hoo.

OpenStudy (anonymous):

Its B an identity is an equation that is true for all real values in its domain